# CitDet: A Benchmark Dataset for Citrus Fruit Detection

### v2

@@ I Introduction
Figure 1: The CitDet dataset contains precise bounding box object annotations for fruit on tree and fruit on ground (top row). It also has images from multiple different tree rows including a large variety of citrus (bottom row).
Fruit detection and counting in orchards are crucial tasks for agricultural automation. They can be used to reduce routine farming and breeding activities as well as provide insightful estimates for harvest and forthcoming growing seasons. Moreover, accurate fruit detection enables the possibility of robotic harvesting, which has the potential to eliminate one of the most labor-intensive processes for growers. Many imaging and sensing technologies have been used for detecting fruit such as hyperspectral [1], laser scanning [2], thermal [3], and red-green-blue depth (RGBD) sensors [4, 5], yet the most common technology is the standard RGB camera.
Although conventional RGB cameras are widely accessible, they present several challenges for in-orchard fruit detection such as variation in appearance, irregular lighting, and severe occlusion. Recent works have used deep learning to overcome these difficulties [6]. However, due to the lack of standardized benchmark datasets for agricultural automation, it is difficult to compare these methods with each other. Citrus datasets (e.g., CitrusFarm [7]) have made large amounts of in-orchard citrus images available, yet there still remains a lack of annotated citrus for detection. To tackle this problem we establish a new benchmark dataset, CitDet [8], for citrus fruit detection and counting in orchard settings together with a comprehensive analysis of state-of-the-art object detection algorithms, Fig. 1.
Computer vision techniques combined with deep learning are appealing in agricultural automation due to their powerful prediction capabilities and non-invasive nature. Nonetheless, such methods require huge amounts of data to perform with high accuracy. While there exists large datasets (e.g., COCO [9]) that have allowed for the development of new algorithms, many automation tasks require custom datasets to achieve meaningful results. For example, even though COCO contains a class for oranges, an orange detector trained solely from the dataset instances will perform poorly in an orchard setting. This is due to the complex background scenes in COCO, which require the detector to have more background classification machinery resulting in less orange classification mechanisms [10]. Another reason for the weak performance stems from many instances of oranges with very little or no occlusions in COCO. In orchard fields, oranges can be occluded by other instances of oranges as well as the surrounding foliage. Furthermore, orange instances will not be a large percentage of the image when an entire tree is imaged. These pitfalls highlight the need for unique datasets to carry out fruit detection, yield mapping, and much more.
@@ I-A Citrus Greening
Citrus greening, also known as Huanglongbing (HLB), is a serious bacterial disease that affects citrus trees and is threatening the existence of the citrus industry, particularly in the southeastern United States [11]. HLB is transmitted by the Asian citrus psyllid, a small insect that feeds on the leaves and shoots of citrus trees. Symptoms of citrus greening include yellowing of the leaves, uneven ripening of the fruit, and reduced fruit size and yield, both in production of fruit as well as premature fruit loss with fruits dropping to the ground. If left untreated, the disease will ultimately lead to the death of the tree. Currently, there are no effective treatments to cure trees affected by citrus greening. The best way to control the spread of the disease is through a combination of methods such as cultural, biological, chemical, and genetic control along with nutritional management. Nevertheless, none of these countermeasures are foolproof. HLB continues to spread and cause major economic losses for citrus growers.
@@ I-B Challenges in Detecting Citrus Fruits in Native Orchard Environments
While public datasets for citrus detection exist (e.g., OrangeSORT [12]), to the best of our knowledge CitDet is the first dataset to include images of citrus trees affected by HLB. As fruit from an infected tree are likely to be discolored and deformed, labeled examples of affected fruits enhance the robustness of citrus detectors in orchard environments plagued by the disease. Images in contemporary datasets are also low resolution, thus limiting the amount of information to learn and predict from. Finally, the aforementioned datasets only annotate fruit on trees. This can be an issue when training a detector for yield estimation since the model will not learn the difference between dropped fruit and fruit still on the tree, which can be helpful in identifying symptoms of HLB or secondary fungus associated with post bloom fruit drop. In contrast, CitDet aims to enhance automation for citrus orchards through high-resolution images of HLB infected citrus trees with accurate annotations for both fruit on tree and fruit on ground. Our contributions are as follows.
@@
The remainder of this paper is organized as follows. Relevant related literature is reviewed in Section II. We discuss the image collection and annotation procedures in Section III and Section IV, respectively. Dataset statistics are outlined in Section V and an evaluation of the benchmark algorithms is presented in Section VI. In Section VII, we conclude and mention future work.
@@ II Related Work
Benchmark datasets provide the means to train and evaluate new algorithms, they are essential to the advancement of deep learning methods. Datasets such as ImageNet [13], Pascal VOC [14], and COCO [9], have publicly released millions of labeled images that contain a variety of classes and objects. Even though these datasets have enabled breakthroughs in image classification, object detection, and object segmentation, there is still a high demand for specialized datasets in agricultural automation. While the number of agricultural datasets continues to grow, there still remains a lack of data on fruit affected by common diseases.
@@ II-A Fruit Detection
Many works use the YOLO [23] family of object detectors for fruit tracking and yield estimation (e.g., [24, 25, 12]). For instance, Wu et al. [26] modified the YOLOv5 [27] loss function to detect banana bunches at harvest time via extended intersection over union [28]. In addition, DeeplabV3+ [29] was utilized to segment banana bunches for counting during sterile bud removal time. Faster R-CNN [30], and the instance segmentation Mask R-CNN [31] variant, have also been successfully employed. Chu et al. [32] added a feature suppression network to the end of Mask R-CNN to filter unwanted foliage from the masks. Even though deep learning is advantageous for fruit detection, the task remains challenging due to the unstructured background prevalent in orchards. Tang et al. [33] provided an extensive review on strategies involving sensor, image processing, and algorithm optimization for orchard environments.
In-orchard detection of fruits is vital for agricultural tasks including fruit yield estimation and automated picking. For example, Dorj et al. [34] leveraged image processing, and the LAB color space, to detect oranges and estimate yield when compared against human observation. Faster R-CNN was used to detect citrus for automated size estimation using drones [35]. Behera et al. [36] also utilized Faster R-CNN for detection and yield estimation of various fruits. YOLOv4 [37] was deployed to detect apples and the trunks of trees by Gao et al. [38]. The detections are used downstream for apple counting by tracking of the tree trunks in videos. The estimation of tree-level yield by Vijayakumar et al. [39] involved the integration of citrus fruit detections from YOLOv3, hyperspectral data, and tree-level traits as inputs for classical machine learning algorithms. Wang et al. [40] used YOLOv5 to detect individual litchi and Mask R-CNN to segment the fruit and determine point locations for picking.
TABLE I: Comparison of fruit detection datasets
Fruit Detection Fruit # Images # Annotations # Classes Resolution Ground Truth Public
Bargoti et al. [16] Apple 841 5765 1 308 x 202 Circles Yes
Stein et al. [41] Mango 1404 7065 1 500 x 500 Circles Yes
Zhang et al. [12] Citrus   -   - 1 800 x 1240 Boxes Yes
Häni et al. [17] Apple 1001 41,325 1 1280 x 720 Polygons Yes
Hou et al. [25] Citrus 4,855 17,567 1 1280 x 720 Boxes Yes
CitDet (ours) Citrus 579 32,502 2 2448 x 3264 Boxes Yes
@@ II-B Comparison of Datasets
Another issue is that many of the datasets, for example the MinneApple [17] dataset, contain fruit on the ground without any labels. This can lead to ambiguity when training a model as it must learn to classify fruit on the ground as “no object” even though the features are similar to the features of fruit on the tree. Our dataset aims to fix these issues and provide researchers with a means to evaluate and compare their algorithms. To the best of our knowledge, CitDet is the first fruit detection dataset to provide full-resolution images with annotations for both fruit on trees and on the ground. It is the only citrus dataset to contain images of entire citrus trees. In addition, we include a variety of citrus tree species imaged at different stages of maturity under varying illumination conditions to avoid overfitting.
@@ III Image Collection
Data was collected over the course of one year to allow the fruit to be imaged at different stages of the ripening cycle during citrus production from October to March. Due to the nature of HLB and the mixture of fruit maturity, CitDet contains fruit of different colors and sizes. All trees were imaged from both the sunny and shady sides of the tree row and imaging was done over multiple days to account for variations in weather and lighting conditions. The imaged trees consist of over 50 different varieties of citrus. All images include a timestamp, tree ID, and the side of the tree imaged in the file name.
@@ IV Image Annotation
To construct CitDet, we followed standard annotation methods found in other benchmark object detection datasets. The Roboflow [43] annotation tool was used to label the image data with bounding boxes. Specifically, we labeled all visible fruit in the foreground regardless of occlusion, while fruit in the background was only labeled for instances with no occlusion. We do not provide any annotations for fruits that are fully occluded and thus not visible. Each of the labeled objects in an image are classified as either fruit on tree or fruit on ground.
CitDet is unique for fruit detection in that we provide labels for fruit that has already fallen. This allows detection methods to accurately identify fruit that contributes to the overall fruit yield. Since a single image can contain a large number of object instances and heavy amounts of occlusion, labeling images is a time-consuming task. For instance, annotating a single image can take up to half an hour. Therefore, the task of instance labeling was split among several internal volunteers and each image was assigned to only one volunteer for annotation. After labeling, each image was then reviewed by a single expert to validate the quality of the annotations.
@@ V Dataset Statistics
[…]
CitDet also provides a challenge for state-of-the-art detection algorithms due to the small average size of the object instances. Small objects are generally harder to detect and require specialized network structures. In ImageNet, Pascal VOC, and COCO, only 50% of the objects occupy less than 10% of an image while the remaining 50% of the objects occupy between 10% and 100% of an image. Conversely, over 90% of the objects in CitDet take up less than 10% of an image. The average object in our dataset is only $50\times 50$ pixels (Fig. 2), which corresponds to less than 0.1% of an image. In contrast, objects in MinneApple occupy around 0.17% of an image. Another key difference between CitDet and MinneApple is the format of the object annotations, bounding boxes versus polygons, respectively. We provide annotations in bounding box format allowing for maskless frameworks to be easily applied. This eliminates the need to infer bounding boxes from masks, which can induce errors due to the highly cluttered and occluded nature of in-orchard fruit.
@@ VI Algorithmic Analysis
We applied multiple state-of-the-art object detection algorithms on CitDet to establish a baseline for future work. For all experiments, we used the same 460 images from our dataset for training while the remaining 119 images were utilized as the test set. We divided the experiments into two tasks: (i) detection on the whole (full-resolution) image dataset, and (ii) detection on the tiled image dataset. The tiled image dataset was created by splitting every image into 9 separate images using a $3\times 3$ cropping method. For any bounding box split by the cropping, a bounding box for each portion of the object was created in each of the new images resulting in slightly more object instances. The resulting dataset contains 5,211 images at a resolution of $816\times 1088$.
@@ VI-A Fruit Detection
In all experiments, the models were trained using an NVIDIA Tesla T4 GPU and restricted to similar levels of compute power. Every algorithm tested used pretrained COCO weights for initialization and they were fine-tuned for 50 epochs on the tiled dataset. The models tested on the whole image dataset used pretrained weights from the tiled dataset and were fine-tuned for 50 epochs on resized images of $704\times 704$ resolution. We evaluated each of the algorithms on the test set and reported the average performance for each class as well as the mean average performance of all classes.
[…]
@@
[…]
Faster R-CNN 41.3 M 22.0 51.5 14.0 21.8 44.2   -
YOLOv5 21.1 M 34.8 70.0 29.8 9.0 36.9 46.3
YOLOv7 36.5 M 40.6 77.9 36.7 12.3 42.6 49.7
@@
[…]
YOLOS 30.9 M 32.4 70.7 24.6 10.3 35.8 47.9
DETR 41.3 M 35.0 72.8 28.8 9.8 38.7 49.8
Faster R-CNN 41.3 M 37.2 76.0 31.5 24.1 44.4 45.9
YOLOv5 21.1 M 44.9 81.9 43.4 18.2 49.2 50.6
YOLOv7 36.5 M 45.5 83.1 43.9 19.2 49.4 52.1
YOLOv7: YOLOv7 [47] improves upon real-time object detection performance by designing a trainable “bag-of-freebies” set of methods. The detector addresses two new issues in object detection evolution by proposing the “extend” and “compound scaling” techniques. The proposed methods reduce the number of parameters and amount of computation by approximately 40% and 50%, respectively. They also result in faster inference speed and higher detection accuracy.
DETR: The DETR [48] model for object detection was introduced by Facebook Artificial Intelligence Research. It uses a ResNet-50 backbone to embed image features as input to a transformer encoder-decoder. The architecture utilizes a standard transformer encoder-decoder with the addition of object query tokens passed to the decoder. The network is trained using a bipartite matching loss thus removing the need for non-maximum suppression. We used sine positional embeddings for both the encoder and decoder.

### v3

@@ I Introduction
Figure 1: The CitDet dataset contains precise bounding box object annotations for fruit on tree (purple) and fruit on ground (yellow) (top row). It also has images from multiple different tree rows including a large variety of citrus (bottom row).
Fruit detection and counting in orchards are crucial tasks for agricultural automation. They can be used to reduce routine farming and breeding activities as well as provide insightful estimates for harvest and forthcoming growing seasons. Moreover, accurate fruit detection enables the possibility of robotic harvesting, which has the potential to eliminate one of the most labor-intensive processes for growers. Many imaging and sensing technologies have been used for detecting fruit such as hyperspectral [1], laser scanning [2], thermal [3], and RGBD sensors [4, 5], yet the most common technology is the standard RGB camera.
Although conventional RGB cameras are widely accessible, they present several challenges for in-orchard fruit detection such as variation in appearance, irregular lighting, and severe occlusion. Recent works have used deep learning to overcome these difficulties [6]. However, due to the lack of standardized benchmark datasets for agricultural automation, it is difficult to compare these methods with each other. Citrus datasets (e.g., CitrusFarm [7]) have made large amounts of in-orchard citrus images available, yet there still remains a lack of annotated citrus for detection. To tackle this problem we establish a new benchmark dataset, CitDet [8, 9], for citrus fruit detection and counting in orchard settings together with a comprehensive analysis of state-of-the-art object detection algorithms, Fig. 1.
Computer vision techniques combined with deep learning are appealing in agricultural automation due to their powerful prediction capabilities and non-invasive nature. Nonetheless, such methods require huge amounts of data to perform with high accuracy. While there exists large datasets (e.g., COCO [10]) that have allowed for the development of new algorithms, many automation tasks require custom datasets to achieve meaningful results. For example, even though COCO contains a class for oranges, an orange detector trained solely from the dataset instances will perform poorly in an orchard setting. This is due to the complex background scenes in COCO, which requires the detector to have more background classification machinery resulting in less orange classification mechanisms [11]. Another reason for the weak performance stems from many instances of oranges with very little or no occlusions in COCO. In orchard fields, oranges can be occluded by other instances of oranges as well as the surrounding foliage. Furthermore, orange instances will not be a large percentage of the image when an entire tree is imaged. These pitfalls highlight the need for unique datasets to carry out fruit detection, yield mapping, and much more.
@@ I-A Citrus Greening
Citrus greening, also known as Huanglongbing (HLB), is a serious bacterial disease that affects citrus trees and is threatening the existence of the citrus industry, particularly in the southeastern United States [12]. HLB is transmitted by the Asian citrus psyllid, a small insect that feeds on the leaves and shoots of citrus trees. Symptoms of citrus greening include yellowing of the leaves, uneven ripening of the fruit, and reduced fruit size and yield, both in the production of fruit as well as premature fruit loss with fruits dropping to the ground. If left untreated, the disease will ultimately lead to the death of the tree. Currently, there are no effective treatments to cure trees affected by citrus greening. The best way to control the spread of the disease is through a combination of methods such as cultural, biological, chemical, and genetic control along with nutritional management. Nevertheless, none of these countermeasures are foolproof. HLB continues to spread and cause major economic losses for citrus growers.
@@ I-B Challenges in Detecting Citrus Fruits in Native Orchard Environments
While public datasets for citrus detection exist (e.g., OrangeSORT [13]), to the best of our knowledge CitDet is the first dataset to include images of citrus trees affected by HLB, accompanied by paired ground-truth metadata for fruit count. As fruit from an infected tree are likely to be discolored and deformed, labeled examples of affected fruits enhance the robustness of citrus detectors in orchard environments plagued by the disease. Images in contemporary datasets are also low resolution, thus limiting the amount of information to learn and predict from. Finally, the aforementioned datasets only annotate fruit on trees. This can be an issue when training a detector for yield estimation since the model will not learn the difference between dropped fruit and fruit still on the tree, which can be helpful in identifying symptoms of HLB or secondary fungus associated with post bloom fruit drop. In contrast, CitDet aims to boost automation for citrus orchards through high-resolution images of HLB infected citrus trees with accurate annotations for both fruit on tree and fruit on ground. We established an interdisciplinary team to address these challenges, including subject matter experts in the biological system. Our contributions are as follows.
@@
  • We make these images, annotations, and associated fruit count metadata available in multiple public spaces to best serve the broader research communities. The CitDet dataset is available through MavMatrix [8] and the USDA Ag Data Commons [9].
@@ II Related Work
Benchmark datasets provide the means to train and evaluate new algorithms, they are essential to the advancement of deep learning methods. Datasets such as ImageNet [14], Pascal VOC [15], and COCO [10], have publicly released millions of labeled images that contain a variety of classes and objects. Even though these datasets have enabled breakthroughs in image classification, object detection, and object segmentation, there is still a high demand for specialized datasets in agricultural automation. Despite the fact that the number of agricultural datasets continues to grow, there remains a lack of data on fruit affected by common diseases.
TABLE I: Comparison of fruit detection datasets
Fruit Detection Fruit # of Images # of Annotations # of Classes Resolution Ground Truth
Bargoti et al. [16] Apple 841 5,765 1 308 $\times$ 202 Circles
Stein et al. [17] Mango 1,404 7,065 1 500 $\times$ 500 Circles
Häni et al. [18] Apple 1,001 41,325 1 720 $\times$ 1,280 Polygons
Hou et al. [19] Citrus 4,855 17,567 1 1,280 $\times$ 720 Boxes
CitDet Citrus 579 44,233 2 2,448 $\times$ 3264 Boxes
CitDet (Tiled) Citrus 5,211 46,630 2 816 $\times$ 1,088 Boxes
@@ II-A Fruit Detection
Many works use the YOLO [26] family of object detectors for fruit tracking and yield estimation (e.g., [27, 19, 13]). For instance, Wu et al. [28] modified the YOLOv5 [29] loss function to detect banana bunches at harvest time via extended intersection over union (IoU) [30]. In addition, DeepLabV3+ [31] was utilized to segment banana bunches for counting during sterile bud removal time. Faster R-CNN [32], and the instance segmentation variant Mask R-CNN [33], have also been successfully employed. Chu et al. [34] added a feature suppression network to the end of Mask R-CNN to filter unwanted foliage from the masks. Even though deep learning is advantageous for fruit detection, the task remains challenging due to the unstructured background prevalent in orchards. Tang et al. [35] provided an extensive review on strategies involving sensor, image processing, and algorithm optimization for orchard environments.
In-orchard detection of fruits is vital for agricultural tasks including fruit yield estimation and automated picking. For example, Dorj et al. [36] leveraged image processing, and the LAB color space, to detect oranges and estimate yield when compared against human observation. Faster R-CNN was used to detect citrus for automated size estimation using drones [37]. Behera et al. [38] also utilized Faster R-CNN for detection and yield estimation of various fruits. YOLOv4 [39] was deployed to detect apples and the trunks of trees by Gao et al. [40]. The detections are used downstream for apple counting by tracking of the tree trunks in videos. The estimation of tree-level yield by Vijayakumar et al. [41] involved the integration of citrus fruit detections from YOLOv3, hyperspectral data, and tree-level traits as inputs for classical machine learning algorithms. Wang et al. [42] used YOLOv5 to detect individual lychee and Mask R-CNN to segment the fruit and determine point locations for picking.
@@ II-B Comparison of Datasets
Another issue is that many of the datasets (e.g., MinneApple [18]) contain fruit on the ground without any labels. This can lead to ambiguity when training a model as it must learn not to detect the fruit on the ground, which can have similar features to that of the fruit on the tree. Other works attempt to solve this issue by treating the unlabeled fruit as a “no object” class and through the introduction of new loss terms to penalize false positive predictions [19]. Our dataset aims to fix these issues and provide researchers with a means to evaluate and compare their algorithms. To the best of our knowledge, CitDet is the first fruit detection dataset to provide full-resolution images with annotations for both fruit on trees and on the ground. It is the only citrus dataset to contain images of entire citrus trees. In addition, we include a variety of citrus tree species imaged at different stages of maturity under varying illumination conditions to avoid overfitting.
@@ III Image Collection
Data was collected over the course of one year to allow the fruit to be imaged at different stages of the ripening cycle during citrus production from October to March. Due to the nature of HLB and the mixture of fruit maturity, CitDet contains fruit of different colors and sizes. All trees were imaged from both the sunny and shady sides of the tree row and imaging was done over multiple days to account for variations in lighting and weather conditions. The imaged trees consist of 60 different varieties of citrus, representing 8 different species of citrus and other related genera. All images include a timestamp, tree ID, and the side of the tree imaged in the file name.
@@ IV Image Annotation
To construct CitDet, we followed standard annotation methods found in other benchmark object detection datasets. The Roboflow [44] annotation tool was used to label the image data with bounding boxes. Specifically, we labeled all visible fruit in the foreground regardless of occlusions, while fruit in the background was only labeled for instances with no occlusions. We do not provide any annotations for fruits that are fully occluded and thus not visible. Each of the labeled objects in an image are classified as either fruit on tree or fruit on ground.
CitDet is unique for fruit detection in that we provide labels for fruit that has already fallen. This allows detection methods to accurately identify fruit that contributes to the overall fruit yield. Since a single image can contain a large number of object instances and a heavy amount of occlusions, labeling images is a time-consuming task. For instance, annotating a single image can take up to half an hour. Therefore, the task of instance labeling was split among several internal volunteers and each image was assigned to only one volunteer to annotate. After labeling, each image was then reviewed by a single expert to validate the quality of the annotations.
@@ V Dataset Statistics
Figure 2: The area distribution of the objects in CitDet. The dataset is mainly comprised of small object instances with an area of less than 502 pixels.
@@
Figure 3: A comparison of the relative instance size (square root of instance area divided by total image area) between fruit detection datasets.
Figure 4: The distribution of object centers in normalized image coordinates for four datasets. CitDet exhibits the greatest spatial diversity among the citrus datasets, with the tiled version exhibiting greater complexity than all the other fruit datasets.
CitDet also provides a challenge for state-of-the-art detection algorithms due to the small average size of the object instances. Small objects are generally harder to detect and require specialized network structures. In ImageNet, Pascal VOC, and COCO, only 50% of the objects occupy less than 10% of an image while the remaining 50% of the objects occupy between 10% and 100% of an image. Conversely, over 90% of the objects in CitDet take up less than 10% of an image. As displayed in Fig. 2, the average object in our dataset is only 50 $\times$ 50 pixels, which corresponds to less than 0.1% of an image. In contrast, objects in MinneApple occupy around 0.17% of an image. Although the average object in CitDet is smaller, the distribution of relative instance size is similar to MinneApple as shown in Fig. 3. The high resolution of CitDet allows for preprocessing techniques (e.g., tiling) to shift the relative distribution and increase spatial complexity. This increases the versatility of the dataset as highlighted in Fig. 4. Another key difference between CitDet and MinneApple is the format of the object annotations, bounding boxes versus polygons, respectively.
@@ VI Algorithmic Analysis
We applied multiple state-of-the-art object detection algorithms on CitDet to establish a baseline for future work. For all the experiments, we used the same 460 images from our dataset for training while the remaining 119 images were utilized as the test set. We divided the experiments into two tasks: (i) detection on the whole (full-resolution) image dataset, and (ii) detection on the tiled image dataset. The tiled image dataset was created by splitting every image into 9 separate images using a 3 $\times$ 3 cropping method. For any bounding box split by the cropping, a bounding box for each portion of the object was created in each of the new images resulting in slightly more object instances. The resulting dataset contains 5,211 images at a resolution of 816 $\times$ 1088.
@@ VI-A Fruit Detection
In all the experiments, the models were trained using an NVIDIA Tesla T4 GPU and restricted to similar levels of compute power. Every algorithm tested used pretrained COCO weights for initialization and they were fine-tuned for 50 epochs on the tiled dataset. The models tested on the whole image dataset used pretrained weights from the tiled dataset and they were fine-tuned for 50 epochs on resized images of 704 $\times$ 704 resolution. We evaluated each of the algorithms on the test set and reported the average performance for each class as well as the mean average performance of all classes.
Detection evaluation metrics: We followed the established evaluation protocols used by other benchmark object detection datasets (e.g., [15, 10, 18]). For a single class, we reported the average precision (AP) as the main evaluation metric. We also recorded the mean AP (mAP) of all classes. For the AP of a class we used an IoU threshold of 0.5, increased it in intervals of 0.05 up to 0.95, and reported the mean of the AP at each threshold. The AP at thresholds of 0.5 and 0.75, as well as the AP for small (APS), medium (APM), and large (APL) objects, are also provided.
@@
Model # of Params AP AP50 APS APM APL
Faster R-CNN 41.3 M 22.0 51.5 21.8 44.2 -
YOLOv5 21.1 M 34.8 70.0 9.0 36.9 46.3
YOLOv7 36.5 M 40.6 77.9 12.3 42.6 49.7
@@
Model # of Params AP AP50 APS APM APL
YOLOS 30.9 M 32.4 70.7 10.3 35.8 47.9
DETR 41.3 M 35.0 72.8 9.8 38.7 49.8
Faster R-CNN 41.3 M 37.2 76.0 24.1 44.4 45.9
YOLOv5 21.1 M 44.9 81.9 18.2 49.2 50.6
YOLOv7 36.5 M 45.5 83.1 19.2 49.4 52.1
YOLOv7: YOLOv7 [48] improves upon real-time object detection performance by designing a trainable “bag-of-freebies” set of methods. The detector addresses two new issues in object detection evolution by proposing the “extend” and “compound scaling” techniques. The proposed methods reduce the number of parameters and the amount of computation by approximately 40% and 50%, respectively. They also result in faster inference speed and higher detection accuracy.
DETR: The DETR [49] model for object detection was introduced by Facebook Artificial Intelligence Research. It uses a ResNet-50 backbone to embed image features as input to a transformer encoder-decoder. The architecture utilizes a standard transformer encoder-decoder with the addition of object query tokens passed to the decoder. The network is trained using a bipartite matching loss, thus removing the need for non-maximum suppression. We used sine positional embeddings for both the encoder and decoder.